OpenStudy (anonymous):

Limit as x approaches 1 (Square root of x+3)-2 / (x-1) Help?

hartnn (hartnn):

did you try to rationalize the numerator ?

hartnn (hartnn):

multiply the numerator and denominator by \(\sqrt{x+3}+2\)

OpenStudy (anonymous):

Yes. That's where I got stuck on @hartnn

hartnn (hartnn):

ok,then lets take the numerator \( .\\ (\sqrt{x+3}-2)(\sqrt{x+3}+2) =.... ? \\~ use, (a+b)(a-b)=a^2-b^2\)

hartnn (hartnn):

treat a as \(\sqrt{x+3} \) and b=2

OpenStudy (anonymous):

Will it be x-1? @hartnn

hartnn (hartnn):

in the numerator, right ?yes!

hartnn (hartnn):

so that gets cancelled with the x-1 in the denominator! what remains ?

OpenStudy (anonymous):

1/(square root of x+3)+2?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

correct! you can just plut x= 1 in that to get the limit! :)

OpenStudy (anonymous):

Can you help me with another question @hartnn ?????

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Thank you !(: Limit as x approaches 0 from the left side Square root of -x @hartnn

hartnn (hartnn):

you can just plug in x =0 in that

OpenStudy (anonymous):

It will be square root of -0..

hartnn (hartnn):

-0 is 0 only square root of 0 is 0 only

OpenStudy (anonymous):

Okay. :) thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Wait @hartnn , help with my last one? :(

hartnn (hartnn):

sure! :)

OpenStudy (anonymous):

Okay. (: Limit as x approaches 3 to the right (Square root of x-3)/(square root of x-1)

hartnn (hartnn):

you can again simply put x=3 because you're not getting any indeterminate form 0/0 or infinity/infinity here

OpenStudy (anonymous):

So it would be correct if it's 0/(square root of 2) ??

hartnn (hartnn):

which = 0 , yes!

OpenStudy (anonymous):

Thank you so much ! Your a life saver :) @hartnn

hartnn (hartnn):

you're most welcome ^_^